The neighborhood store sold 1462 bottles of soft drinks last month. Twice as many bottles of root beer were sold than lemon-lime soda, and twice as many bottles of cola were sold than root beer. How many bottles of each type of soft drink were sold?
Answer by mananth(16946)

Lime soda = x
rootbeer = 2x
cola = 2*2x=4x
...
This is the set up. I assume it is 1463 for round answers
x+2x+4x=1463
7x= 1463
x= 1463/7
x= 209 lime sods
2x = 2*209 = 418 root beer
4x = 4*209 = 836 cola
209+418+836= 1463.
If the total is 1462 then you work it out